OpenStudy (anonymous):

2x – y = -4 -3x + y = -9 show steps please

OpenStudy (anonymous):

2x – y = -4 -3x + y = -9 Add the two. -x=-13 -> x=13 Plug in x in one of the equations and you will find the value of y.

OpenStudy (anonymous):

thank you thank you